# Learning A Physical-aware Diffusion Model Based on Transformer for Underwater Image Enhancement

### v2

@@ 3.2. Physics Prior Generation
Finally, the overall loss $\mathcal{L}_{ppg}$ of PPG branch can be expressed as:
(12) $\mathcal{L}_{ppg}=\lambda_{1}\mathcal{L}_{p}+\lambda_{2}\mathcal{L}_{per}.$
Figure 3. The detailed structure of the proposed physics perception unit (PPU).
3.3. Implicit Neural Reconstruction
Inspired by (Quan et al., 2023; Chen et al., 2024; Yang et al., 2023a) based on implicit neural representations that can encode an image as a continuous function and learn a more robust feature representation, we further introduce Implicit Neural Representations (INRs) to learn robust degradation representations from diverse underwater image, aiming to reduce the difficulty of restoration for diffusion models. INRs trains a multi-layer perceptron (MLP) by learning the following mapping :
(13) $F_{\theta}:\mathbb{R}^{2}\to\mathbb{R}^{3},$
where the dimensions of input refer to the (x, y) spatial coordinates, while the dimensions of output refer to the (R, G, B) color channels. Specifically, following the setting of (Chen et al., 2024; Yang et al., 2023a), underwater image $I$ is transformed to a feature map $\mathbf{E}\in\mathbb{R}^{H\times W\times C}$ with a image resolution of H × W. Meanwhile, the position of each pixel is recorded in a relative coordinate set $\mathbf{X}\in\mathbb{R}^{H\times W\times 2}$, where the value ’2’ means horizontal and vertical coordinates. As shown in Figure 1, we fuse $X$ and $E$, and use $F_{\theta}$ to output image $\boldsymbol{I}_{{\mathrm{inr}}}$, which is expressed as:
(14) $\boldsymbol{I}_{{\mathrm{inr}}}[x,y]=\boldsymbol{F}_{\theta}(\mathbf{E}[x,y],\mathbf{X}[x,y]),$
where $[x,y]$ is the location of a pixel and $\boldsymbol{I}_{{\mathrm{inr}}}[x,y]$ is the generated RGB value. Due to the elimination of high-frequency components during rendering (Yang et al., 2023a), we adopt periodic spatial encoding (Lee and Jin, 2022) to project the pixel coordinates X into a higher dimensional space $\mathbb{R}^{2L}$ for better recovering high-frequency details. The encoding procedure is formulated as:
(15) $\mathbf{X}^{\prime}=\gamma(\mathbf{X}),$
(16) $\gamma(c)=(\cdot\cdot\cdot,\sin(2^{i}c),\cos(2^{i}c),\cdot\cdot\cdot),$
where $\gamma()$ represents a spatial encoding function. $i$ values from 00 to $L-1$, $c\in\mathbb{R}^{2L}$ is the coordinate value of $X$, and $L$ is a hyperparameter that determines the dimension value. Given a GT of underwater image $GT$, The reconstruction loss of INR branch can be expressed as:
(17) $\mathcal{L}_{INR}=||\boldsymbol{I}_{\mathrm{inr}}-GT||_{1}.$
@@ 3.3. Implicit Neural Reconstruction
(18) $x_{c}=\boldsymbol{I}_{\mathrm{inr}}+I.$
3.4. Physics-aware Diffusion Transformer
Physics-aware diffusion transformer (PDT) branch aims to exploit the extracted physics prior information from PPG branch for better guiding underwater image restoration. Fig. 1 shows the overall of PDT branch, which is a U-shaped structure with physics-aware diffusion transformer (PDT) blocks and cross attention module (CAM).
Given the output of the INR branch as condition image $x_{c}\in{R}^{H\times W\times 3}$ and the background light $B^{c}\in{R}^{H\times W\times 3}$ as a physical prior condition, we first obtain the noise sample $x_{t}\in{R}^{H\times W\times 3}$ at time step $t$ according to the forward process, and concatenate $x_{t}$, $B^{c}$ and $x_{c}$ at channel dimension to obtain the input of PDT branch. Then, we obtain the embedding features $\mathbf{F}\in{R}^{H\times W\times C}$ through convolution, where C means the number of channel. $\mathbf{F}$ is encoded and decoded by PDT block, which consists of physics-aware self-attention (PA-SA), physics perception unit and gated multi-scale feed-forward network (GM-FFN). Note that, we introduce a physics perception unit (PPU) to further exploit the beneficial feature-level physics information.
Cross Attention Module. CAM aims to obtain a physics-aware interaction embedding, using which guides the diffusion process. We denote $T^{c}$ and $\mathbf{F}$ as the input of CAM. We use different linear projections to construct $Q$ and $K$ in CAM:
(19) $Q=Conv_{1\times 1}(T^{c}),$
(20) $K=Conv_{1\times 1}(\mathbf{F}).$
Similarly, $V$ of the transmission map $T^{c}$ can be obtained:
(21) $V=Conv_{1\times 1}(T^{c}).$
The output feature map of CAM $T_{out}$ can then be obtained from the formula:
(22) $T_{out}=Softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V.$
Physics-aware Self-attention. As is shown in Figure 2 (a), we first embed the time embedding $\mathbf{t}$ into the input features $\mathbf{F}$ in our PA-SA:
(23) $\tilde{\mathbf{F}}=Norm(\mathbf{F})+\mathbf{t},$
where $Norm$ means layer normalization. We first obtain query ($Q_{f}=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$), key ($K=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$), and value ($V=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$) from the embedding input features $\mathbf{F}$. Meanwhile, we generate query ($Q_{t}=\mathbf{W}_{d}\mathbf{W}_{p}\mathbf{T_{out}}$) from the output feature map of CAM, where $\mathbf{W}_{p}$ and $\mathbf{W}_{d}$ respectively denote 1×1 point-wise convolution and 3×3 depth-wise convolution. We employ physics-aware query-driven attention strategy to obtain the final query $Q$ by fusing $Q_{t}$ and $Q_{f}$, which can be expressed as:
(24) $Q=Conv_{1\times 1}(concat(Q_{f},Q_{t})).$
@@ 3.4. Physics-aware Diffusion Transformer
(25) $\hat{\mathbf{F}}=\mathbf{F}+softmax\left(\mathbf{Q}\mathbf{K}^{T}/\alpha\right)\cdot\mathbf{V},$
@@
Physics Perception Unit. Exploring feature-level physics information (Zheng et al., 2023; Dong and Pan, 2020) achieved success in image dehazing tasks. Our proposed physics perception unit (PPU) aims to further exploit the beneficial feature-level physics information for underwater image enhancement, and learn richer physical features. Our proposed PPU is based on the simplified atmospheric model Eq. (8), as shown in Fig. 3. We can reformulate the physics model to represent the clear image J, which can be obtained by:
(26) $J^{c}(\mathbf{x})=I^{c}(\mathbf{x})\frac{1}{T^{c}(\mathbf{x})}+B^{c}(1-\frac{1}{T^{c}(\mathbf{x})}).$
@@
(27) $k\otimes J=k\otimes(I\odot\frac{1}{t})+k\otimes B(1-\frac{1}{t}),$
@@
(28) $\boldsymbol{KJ}=\boldsymbol{K}\boldsymbol{T}_{1}\boldsymbol{I}+\boldsymbol{K}\boldsymbol{T}_{2}\boldsymbol{B},$
where $\boldsymbol{K},\boldsymbol{J},\boldsymbol{I},\boldsymbol{B},\boldsymbol{T}_{1}$ and $\boldsymbol{T}_{2}$ respectively denote the matrix-vector forms of $k,J,I,B,\frac{1}{t}$ and $(1-\frac{1}{t})$. Note that the diagonal vector of the diagonal matrix $\boldsymbol{T}_{1}$ corresponds to the vectorized form of $\frac{1}{t}$. Assuming that $\boldsymbol{K}$ is an invertible matrix-vector, we can obtain
(29) $\boldsymbol{F}_{1}=\boldsymbol{K}\boldsymbol{T}_{1}\boldsymbol{K}^{-1},$
(30) $\boldsymbol{F}_{2}=\boldsymbol{K}\boldsymbol{T}_{2}\boldsymbol{K}^{-1}.$
And then, Eq. (27) can be express as:
(31) $\boldsymbol{KJ}$ $=(\boldsymbol{K}\boldsymbol{T}_{1}\boldsymbol{K}^{-1})\boldsymbol{K}\boldsymbol{I}+(\boldsymbol{K}\boldsymbol{T}_{2}\boldsymbol{K}^{-1})\boldsymbol{K}\boldsymbol{B}$
$=\boldsymbol{F}_{1}(\boldsymbol{K}\boldsymbol{I})+\boldsymbol{F}_{2}(\boldsymbol{K}\boldsymbol{B}).$
@@
(32) $\tilde{\boldsymbol{J}}=\tilde{\boldsymbol{t}_{1}}\odot\tilde{\boldsymbol{I}}+\tilde{\boldsymbol{t}_{2}}\odot\tilde{\boldsymbol{B}}.$
@@
(33) $\tilde{\boldsymbol{B}}=H(\sigma(\mathrm{Conv}(\mathrm{ReLU}(\mathrm{Conv}(\mathrm{GAP}(\tilde{\boldsymbol{I}}))))))))),$
@@
As the transmission map is non-homogeneous, we cannot apply GAP for the approximations of $\tilde{\boldsymbol{t}_{1}}$ and $\tilde{\boldsymbol{t}_{2}}$. We use different projection networks to obtain $\tilde{\boldsymbol{t}_{1}}$ and $\tilde{\boldsymbol{t}_{2}}$ :
(34) $\tilde{\boldsymbol{t}_{1}}=\sigma(\mathrm{Conv}(\mathrm{ReLU}(\mathrm{Conv}(\mathrm{Conv_{s}}(\tilde{\boldsymbol{I}})))))),$
(35) $\tilde{\boldsymbol{t}_{2}}=\sigma(\mathrm{Conv}(\mathrm{ReLU}(\mathrm{Conv}(\mathrm{Conv_{s}}(\tilde{\boldsymbol{I}})))))).$
where s represents the shared Conv layer. Given $\hat{\mathbf{F}}$ as the input of PPU, we use residual connection to obtain the final output of PPU, which can be expressed as :
(36) $\hat{\mathbf{F_{ppu}}}=PPU(\hat{\mathbf{F}})+\hat{\mathbf{F}}.$
Gated Multi-scale Feed-Forward Network. Finaly, we design a GM-FFN for local feature aggregation. As is shown in Figure 2 (b), we employ GD-FFN-like structure (Zamir et al., 2022) to construct our feed-forward network. In order to expand the receptive field, we employ multi-scale kernel depth-wise convolutions. GM-FFN adopts GELU to ensure the flexibility of feature aggregation.
@@
Figure 4. The visual results of the background scattered light $B^{c}$, the medium transmission map $T^{c}(x)$, and the implicit neural output $I_{inr}$.
@@
Table 1. Quantitative comparison on the UIEBD and LSUI datasets. The best results are highlighted in bold and the second best results are underlined.
Methods UIEWD UWCNN UIECˆ2-Net Water-Net U-color U-shape DM-water Ours

### v3

@@ III-B Physics Prior Generation
Finally, the overall loss $\mathcal{L}_{ppg}$ of the PPG branch can be expressed as:
$\mathcal{L}_{ppg}=\lambda_{1}\mathcal{L}_{p}+\lambda_{2}\mathcal{L}_{per}.$ (13)
Figure 3: The detailed architecture of our designed (a) physics-aware self-attention (PA-SA) and (b) gated multi-scale feed-forward network (GM-FFN).
Figure 4: The detailed structure of the proposed physics perception unit (PPU). To the best of our knowledge, it is the first method which is based on the interpretable physics model in the feature space for challenging underwater image enhancement tasks.
III-C Implicit Neural Reconstruction
Inspired by [37, 3, 52] based on implicit neural representations that can encode an image as a continuous function and learn a more robust feature representation, we further introduce Implicit Neural Representations (INRs) to learn robust degradation representations from diverse underwater images, aiming to reduce the difficulty of restoration for diffusion models. INRs train a multi-layer perceptron (MLP) by learning the following mapping :
$F_{\theta}:\mathbb{R}^{2}\to\mathbb{R}^{3},$ (14)
$\mathbb{R}^{2}$ denotes the continuous 2D spatial coordinate domain $(x,y)$ and $\mathbb{R}^{3}$ denotes the corresponding RGB color space, i.e., each spatial location is mapped to its RGB color value. Specifically, following the setting of [3, 52], an underwater image $I$ is transformed to a feature map $\mathbf{E}\in\mathbb{R}^{H\times W\times C}$ with an image resolution of H × W. Meanwhile, the position of each pixel is recorded in a relative coordinate set $\mathbf{X}\in\mathbb{R}^{H\times W\times 2}$, where the value ’2’ means horizontal and vertical coordinates. As shown in Figure 2, we fuse $X$ and $E$, and use $F_{\theta}$ to output image $\boldsymbol{I}_{{\mathrm{inr}}}$, which is expressed as:
$\boldsymbol{I}_{{\mathrm{inr}}}[x,y]=\boldsymbol{F}_{\theta}(\mathbf{E}[x,y],\mathbf{X}[x,y]),$ (15)
where $[x,y]$ is the location of a pixel and $\boldsymbol{I}_{{\mathrm{inr}}}[x,y]$ is the generated RGB value. Due to the elimination of high-frequency components during rendering [52], we adopt periodic spatial encoding [19] to project the pixel coordinates X into a higher dimensional space $\mathbb{R}^{2L}$ for better recovering high-frequency details. The encoding procedure is formulated as:
$\mathbf{X}^{\prime}=\gamma(\mathbf{X}),$ (16)
$\gamma(x)=(\cdot\cdot\cdot,\sin(2^{i}x),\cos(2^{i}x),\cdot\cdot\cdot),$ (17)
where $\gamma()$ represents a spatial encoding function. $i$ values from 0 to L1L-1, and $L$ is a hyperparameter that determines the dimension value. Given a GT of underwater image $GT$, The reconstruction loss of INR branch can be expressed as:
[…]
@@ III-C Implicit Neural Reconstruction
$x_{c}=\boldsymbol{I}_{\mathrm{inr}}+I.$ (19)
III-D Physics-aware Diffusion Transformer
Physics-aware diffusion transformer (PDT) branch aims to exploit the extracted physics prior information from PPG branch for better guiding underwater image restoration. Figure 5 showcases the prior results provided by our PPG branch and INR branch. Figure 2 shows the overall of PDT branch, which is a U-shaped structure with physics-aware diffusion transformer (PDT) blocks and cross attention modules (CAM). Given the output of the INR branch as the condition image $x_{c}\in{R}^{H\times W\times 3}$ and the background light $B^{c}\in{R}^{H\times W\times 3}$ as a physical prior condition, we first obtain the noise sample $x_{t}\in{R}^{H\times W\times 3}$ at time step $t$ according to the forward process, and concatenate $x_{t}$, $B^{c}$ and $x_{c}$ at channel dimension to obtain the input of PDT branch. Then, we obtain the embedding features $\mathbf{F}\in{R}^{H\times W\times C}$ through convolution, where C means the number of channel. $\mathbf{F}$ is encoded and decoded by PDT block, which consists of physics-aware self-attention (PA-SA), physics perception unit and gated multi-scale feed-forward network (GM-FFN). Note that, we introduce a physics perception unit (PPU) to further exploit the beneficial feature-level physics information.
Cross Attention Module. CAM aims to obtain a physics-aware interaction embedding, using which guides the diffusion process. We denote $T^{c}$ and $\mathbf{F}$ as the input of CAM. We use different linear projections to construct $Q$ and $K$ in CAM: $Q=Conv_{1\times 1}(T^{c}),K=Conv_{1\times 1}(\mathbf{F}).$ Similarly, $V$ of the transmission map $T^{c}$ can be obtained: $V=Conv_{1\times 1}(T^{c}).$ The output feature map of CAM $T_{out}$ can then be obtained from the formula:
$T_{out}=Softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V.$ (20)
Physics-aware Self-attention. As is shown in Figure 3 (a), we first embed the time embedding $\mathbf{t}$ into the input features $\mathbf{F}$ in our PA-SA: $\tilde{\mathbf{F}}=Norm(\mathbf{F})+\mathbf{t}$. We obtain query ($Q_{f}=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$), key ($K=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$), and value ($V=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$) from the embedding input features $\mathbf{F}$. Meanwhile, we generate query ($Q_{t}=\mathbf{W}_{d}\mathbf{W}_{p}\mathbf{T_{out}}$) from the output feature map of CAM, where $\mathbf{W}_{p}$ and $\mathbf{W}_{d}$ respectively denote 1×1 point-wise convolution and 3×3 depth-wise convolution. We employ physics-aware query-driven attention strategy to obtain the final query $Q$ by fusing $Q_{t}$ and $Q_{f}$, which can be expressed as:
$Q=Conv_{1\times 1}(concat(Q_{f},Q_{t})).$ (21)
@@ III-D Physics-aware Diffusion Transformer
$\hat{\mathbf{F}}=\mathbf{F}+softmax\left(\mathbf{Q}\mathbf{K}^{T}/\alpha\right)\cdot\mathbf{V},$ (22)
@@
Physics Perception Unit. Exploring feature-level physics information [70, 8] achieved success in image dehazing tasks. Our proposed physics perception unit (PPU) aims to further exploit the beneficial feature-level physics information for underwater image enhancement, and learn richer physical features. Our proposed PPU is based on the simplified atmospheric model Eq. (5), as shown in Figure 4. We can reformulate the physics model to represent the clear image J, which can be obtained by:
$J^{c}(\mathbf{x})=I^{c}(\mathbf{x})\frac{1}{T^{c}(\mathbf{x})}+B^{c}(1-\frac{1}{T^{c}(\mathbf{x})}).$ (23)
@@
$k\otimes J=k\otimes(I\odot\frac{1}{t})+k\otimes B(1-\frac{1}{t}),$ (24)
@@
$\boldsymbol{KJ}=\boldsymbol{K}\boldsymbol{T}_{1}\boldsymbol{I}+\boldsymbol{K}\boldsymbol{T}_{2}\boldsymbol{B},$ (25)
where $\boldsymbol{K},\boldsymbol{J},\boldsymbol{I},\boldsymbol{B},\boldsymbol{T}_{1}$ and $\boldsymbol{T}_{2}$ respectively denote the matrix-vector forms of $k,J,I,B,\frac{1}{t}$ and $(1-\frac{1}{t})$. Note that the diagonal vector of the diagonal matrix $\boldsymbol{T}_{1}$ corresponds to the vectorized form of $\frac{1}{t}$.
Then, we can decompose the matrix $\boldsymbol{K}\boldsymbol{T}_{1}$ and $\boldsymbol{K}\boldsymbol{T}_{2}$ into product of two matrices $\boldsymbol{F}_{1}\boldsymbol{K}$ and $\boldsymbol{F}_{2}\boldsymbol{K}$. Eq. (19) can be express as:
$\boldsymbol{KJ}$ $=\boldsymbol{F}_{1}(\boldsymbol{K}\boldsymbol{I})+\boldsymbol{F}_{2}(\boldsymbol{K}\boldsymbol{B}).$ (26)
@@
$\tilde{\boldsymbol{J}}=\tilde{\boldsymbol{t}_{1}}\odot\tilde{\boldsymbol{I}}+\tilde{\boldsymbol{t}_{2}}\odot\tilde{\boldsymbol{B}}$. (27)
@@
[…]
@@
As the transmission map is non-homogeneous, we cannot apply GAP for the approximations of $\tilde{\boldsymbol{t}_{1}}$ and $\tilde{\boldsymbol{t}_{2}}$. $\tilde{\boldsymbol{t}_{1}}$ and $\tilde{\boldsymbol{t}_{2}}$ can be obtained by
$))),\tilde{\boldsymbol{t}_{1}},\tilde{\boldsymbol{t}_{2}}=\sigma(\mathrm{Conv}(\mathrm{ReLU}(\mathrm{Conv}(\mathrm{Conv}(\tilde{\boldsymbol{I}}$)))))), (29)
Gated Multi-scale Feed-forward Network. Finally, we design a GM-FFN for local feature aggregation. As is shown in Figure 3 (b), we employ GD-FFN-like structure [57] to construct our feed-forward network. In order to expand the receptive field, we employ multi-scale kernel depth-wise convolutions. GM-FFN adopts GELU to ensure the flexibility of feature aggregation. Discussion on Prior Multi-scale Works. MB-TaylorFormerv2 [17] introduces multi-scale patch embedding in transformer for image restoration, incorporating varying receptive field sizes, flexible receptive field shapes, and multi-level semantic information. ESTINet [58] proposes a feature-level multi-scale fusion module between the encoder and decoder for video deraining. DDMSNet [60] presents a multi-stage coarse-to-fine multi-scale CNN framework for snow removal, achieving progressive restoration. GridFormer [41] adopts a pixel-level multi-scale transformer architecture for image restoration, effectively leveraging complementary multi-scale information in the pixel space. Other works [61, 59, 67, 38, 69] also explore multi-scale information fusion and utilization at both the pixel and feature levels. In contrast, our work investigates multi-scale information utilization within the feed-forward network (FFN) of transformers, enhancing the FFN’s ability to capture multi-scale features.
@@
Figure 5: The visual results of the background scattered light $B^{c}$, the medium transmission map $T^{c}(x)$, and the implicit neural output $I_{inr}$.
Figure 6: Qualitative comparison with other SOTA methods on the LSUI dataset. (a) input, (b) UIEWD,(c) UWCNN, (d) WaterNet, (e) UIEC2Net, (f) Ucolor, (g) GUPDM, (h) CECF, (i) DM-water, (j) HCLR, (k) $S$-UIE, (l) PA-Diff.
Figure 7: Qualitative comparison with other SOTA methods on the UIEB dataset. (a) input, (b) UIEWD,(c) UWCNN, (d) WaterNet, (e) UIEC2Net, (f) Ucolor, (g) GUPDM, (h) CECF, (i) DM-water, (j) HCLR, (k) $S$-UIE, (l) PA-Diff.
TABLE I: Implementation details of our PA-Diff.
Hyperparameter Value
Optimizer Adam
Learning Rate $1\times 10^{-4}$ (fixed)
Batch Size 6
Total Iterations 1,050,000
Input Resolution $256\times 256$
Diffusion Timesteps 2,000
Beta Schedule Linear ($1\times 10^{-6}$ to $1\times 10^{-2}$)
UNet Inner Channel 48
Channel Multiplier $[1,2,4,8,8]$
Attention Resolution 16
Dropout 0.2
Norm Groups 24
GPU Hardware NVIDIA RTX 3090 (24GB)
Training Time per Epoch $\sim$8.6 minutes
Training Time per Iteration $\sim$0.8 seconds
@@
TABLE II: Quantitative comparison on the UIEB and LSUI datasets. The best results are highlighted in bold and the second best results are underlined.
Method UIEB LSUI
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ UCIQE $\uparrow$ UIQM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ UCIQE $\uparrow$ UIQM $\uparrow$
UIEWD [27] 13.68 0.696 0.437 174.129 0.607 4.132 14.45 0.658 0.429 101.326 0.547 4.369
UWCNN [2] 13.94 0.716 0.335 91.048 0.551 4.082 16.59 0.706 0.361 61.202 0.551 4.103
WaterNet [33] 22.76 0.906 0.129 42.432 0.615 4.341 20.78 0.804 0.234 37.900 0.569 4.349
UIEC2Net [45] 20.04 0.818 0.178 49.071 0.606 4.369 24.25 0.854 0.142 37.425 0.591 4.307
Ushape [21] 18.15 0.833 0.210 68.359 0.555 4.235 23.66 0.845 0.158 29.051 0.573 4.351
Ucolor [20] 20.08 0.830 0.184 73.245 0.587 4.236 23.98 0.864 0.166 35.279 0.566 4.235
GUPDM [29] 19.90 0.877 0.157 41.825 0.573 4.226 24.48 0.869 0.119 29.360 0.578 4.380
CECF [7] 22.20 0.841 0.135 32.370 0.619 4.374 26.58 0.886 0.115 19.636 0.590 4.355
DM-water [40] 22.78 0.894 0.113 29.249 0.598 4.217 26.95 0.881 0.143 21.433 0.589 4.257
WF-Diff [64] 22.84 0.898 0.114 28.482 0.609 4.293 26.07 0.884 0.121 20.846 0.587 4.369
HCLR [73] 22.87 0.919 0.112 27.864 0.607 4.292 26.33 0.880 0.108 27.211 0.593 4.408
SS-UIE [32] 21.98 0.907 0.121 36.479 0.589 4.263 24.41 0.840 0.164 26.196 0.575 4.318
PA-Diff 23.47 0.897 0.106 23.631 0.621 4.385 27.25 0.889 0.103 20.463 0.594 4.393
TABLE III: Quantitative comparison on EUVP, U45, and Challenge subsets. The best results are highlighted in bold and the second best results are underlined.
Method EUVP U45 Challenge
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ UCIQE $\uparrow$ UIQM $\uparrow$ UCIQE $\uparrow$ UIQM $\uparrow$ UCIQE $\uparrow$ UIQM $\uparrow$
UIEWD [27] 13.93 0.623 0.452 91.712 0.504 4.063 0.510 4.057 0.550 3.641
UWCNN [2] 17.50 0.692 0.341 55.767 0.570 4.157 0.534 4.075 0.535 3.671
WaterNet [33] 18.43 0.770 0.308 39.888 0.585 4.182 0.552 4.055 0.579 3.839
UIEC2Net [45] 24.04 0.818 0.180 29.755 0.594 4.248 0.575 4.105 0.581 3.935
Ushape [21] 25.53 0.818 0.139 24.914 0.587 4.238 0.567 4.139 0.530 3.820
Ucolor [20] 25.68 0.871 0.159 26.114 0.586 4.240 0.560 4.014 0.542 3.683
GUPDM [29] 25.18 0.848 0.114 20.744 0.590 4.399 0.566 4.144 0.543 3.827
CECF [7] 28.02 0.861 0.126 22.630 0.596 4.252 0.585 4.128 0.575 3.899
DM-water [40] 26.45 0.852 0.183 24.484 0.598 4.266 0.592 4.139 0.586 3.919
WF-Diff [40] 27.24 0.851 0.128 21.517 0.592 4.269 0.584 4.137 0.584 3.926
HCLR [73] 27.08 0.850 0.117 25.114 0.591 4.288 0.584 4.117 0.562 3.792
SS-UIE [32] 26.00 0.824 0.160 24.829 0.587 4.353 0.584 4.105 0.557 3.970
PA-Diff 29.12 0.880 0.101 18.769 0.599 4.273 0.593 4.147 0.586 3.954
Figure 8: Qualitative comparison with other SOTA methods on the Challenge dataset. (a) input, (b) UIEWD,(c) UWCNN, (d) WaterNet, (e) UIEC2Net, (f) Ucolor, (g) GUPDM, (h) CECF, (i) DM-water, (j) HCLR, (k) $S$-UIE, (l) PA-Diff.
Figure 9: Visual analysis with ablation study on the LSUI datasets.
Figure 10: Comparative analysis of diffusion-based underwater image restoration with and without physical priors.
Figure 11: Multi-view visual comparison on UwMVS real-world test set [50].
Figure 12: Limitations. The output images refer to the reconstruction images of PPG branch.
Figure 13: Visual results under extreme low-light conditions.
IV Experiments
IV-A Experimental Settings
Implementation Details. Our network, implemented using PyTorch 1.7, underwent training and testing on an NVIDIA GeForce RTX 3090 GPU. We employed the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate was established at 0.0001. During the training stage, to balance the batch size and image size, the batch size and image resolution are set to 6 and 256 × 256, respectively. The number of training iterations reached one million. The pixel values of the image are normalized to [-1,1]. The time step of the diffusion model is set to 2000. A skip sampling strategy is employed with 10 sampling steps to balance enhancement performance and inference runtime. Table I shows the key details regarding the implementation of our PA-Diff.
Datasets. We utilize the UIEB dataset [21] and the LSUI dataset [33] for training and evaluating our model. The UIEB dataset comprises 890 underwater images with corresponding labels. Out of these, 800 images are allocated for training, and the remaining 90 are designated for testing. LSUI dataset contains 4279 underwater images and their corresponding high-quality images. Compared with UIEB, LSUI contains diverse underwater scenes, object categories and deep-sea and cave images. In the paper, the LSUI dataset is randomly partitioned into 3879 images for training and 400 images for testing. In addition, to verify the generalization, we use the reference based benchmark EUVP [15], which contains 515 paired test samples, along with the non-reference based benchmarks U45 [22] and Challenge [21], which contain 45 and 60 underwater images, respectively.
Comparison Methods. We conduct a comparative analysis with seven SOTA UIE methods, namely UIECˆ2-Net [45], Water-Net [33], UWCNN [2], UIEWD [27], U-color [20], U-shape [21], GUPDM [29], CECF [7], DM-water [40], WF-Diff [64], HCLR [73], and SS-UIE [32]. To ensure a fair and rigorous comparison, we utilize the provided source codes from the respective authors and adhere strictly to the identical experimental settings across all evaluations.
Evaluation Metrics. We employ PSNR and SSIM [46] for quantitative comparisons at both pixel and structural levels. Additionally, LPIPS [62] and FID [13] are utilized to evaluate perceptual performance. For non-reference benchmarks U45 and Challenge, we introduce UIQM [31] and UCIQE [51] to evaluate our method.
IV-B Results and Comparisons
Table II shows the quantitative results compared with different SOTA methods on the UIEB and LSUI datasets. We mainly use PSNR, SSIM, LPIPS, FID, UCIQE and UIQM as our quantitative indices for UIEBD and LSUI datasets. The results show that our algorithm outperforms state-of-the art methods obviously, especially in terms of all metrics. Figure 6 and 7 present visual comparison results with other methods on the LSUI and UIEB datasets, with selected examples randomly chosen from each dataset. Furthermore, Table III presents a comparison using pre-trained models on LSUI across multiple datasets, demonstrating the superior generalization capability of our method. Furthermore, in order to more effectively validate the generalization performance of our method on real-world images, we present enhancement results from the Challenge dataset in Figure 8. Our method consistently generates more natural and visually pleasing results, strongly proving that PA-Diff has good generalization performance for real-world underwater images.
TABLE IV: Ablation with application of diffusion technology.
Methods LSUI EUVP
Name PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/o Diffusion 27.13 0.883 0.127 27.75 0.868 0.153
PA-Diff (Ours) 27.25 0.889 0.103 29.12 0.880 0.101
TABLE V: Ablation study on the LSUI dataset. $T_{c}$ and $B_{c}$ refer to transmission map and background light prior information in the PPG branch, respectively. GBB means Gaussian blur module. INR represents our INR branch. PSE refers to the periodic spatial encoding [19], aiming to better recover high-frequency details. CAM and PPU denote the cross attention module and physics perception unit, respectively. GMFFNGM-FFN means multi-scale operation in the feed-forward network.